# Self-Interacting Dark Matter with Mass Segregation: A Unified Explanation of Dwarf Cores and Small-Scale Lenses

### v1

@@ Abstract
In two-component SIDM models with inter-species interactions, mass segregation arises naturally from collisional relaxation, enhancing central densities and gravothermal evolution without requiring large cross sections. We propose a model with velocity-dependent interactions, both within and between species, that connects observations across several halo mass scales while remaining consistent with cluster-scale constraints. This combination enables modest mass segregation in low-mass and typical-concentration halos, consistent with recent dwarf galaxy clustering measurements. Using cosmological zoom-in simulations and controlled isolated halo studies, we show that this model produces dwarf galaxy cores that grow over time, explains the structure of dark perturbers observed in strong lensing systems, and significantly increases the number and efficiency of small-scale lenses, consistent with the galaxy-galaxy strong lensing excess reported in clusters. Our results establish mass segregation in two-component SIDM as a self-consistent and testable model capable of simultaneously addressing multiple small-scale challenges in structure formation.
@@
Introduction— Observations of large-scale structure have established the standard cosmological model with cold dark matter (CDM). However, on small scales, persistent discrepancies continue to challenge this framework, motivating alternative interpretations Bullock and Boylan-Kolchin (2017); Salucci (2019); Gentile et al. (2004); van Dokkum et al. (2022); Bonaca et al. (2018). Among these, self-interacting dark matter (SIDM) has long been considered a promising candidate Spergel and Steinhardt (2000); Tulin and Yu (2018). Through gravothermal evolution, SIDM can naturally produce both cored and cuspy halo density profiles, providing a flexible mechanism to explain the diversity of galactic rotation curves, the extreme central densities observed in ultra-diffuse and ultra-compact galaxies, and the early formation of supermassive black holes Kaplinghat et al. (2016); Robertson et al. (2019); Despali et al. (2019); Essig et al. (2019); Santos-Santos et al. (2020); Turner et al. (2021); Fischer et al. (2021); Slone et al. (2023); Zeng et al. (2022); Gilman et al. (2021); Correa et al. (2022); Yang and Yu (2022); Adhikari et al. (2022); Outmezguine et al. (2023); Silverman et al. (2022); Yang et al. (2024); Zhong et al. (2023); Zeng et al. (2025); Zhang et al. (2025a, 2024); Ragagnin et al. (2024); Nadler et al. (2025); Kong and Yu (2025); Jiang et al. (2025). It has also been proposed as a solution to the final-parsec problem in binary black hole mergers Alonso-Álvarez et al. (2024). Recently, a striking development by Zhang et al. Zhang et al. (2025b) revealed an unexpected clustering pattern in dwarf galaxies, favoring a self-interaction cross section of $\sigma/m\sim 0.3~{}\rm cm^{2}/g$ in their host halos.
@@
Despite accumulating evidence, a unified SIDM explanation for all small-scale phenomena remains elusive. In particular, the relatively small cross section inferred from dwarf galaxy clustering appears inconsistent with the onset of core collapse required to explain dense subhalos observed via strong lensing, which demands $\sigma/m\gtrsim 100~{}\rm cm^{2}/g$ Nadler et al. (2023); Li et al. (2025). Such compact substructures have been inferred in systems like SDSS J0946+1006 and JVAS B1938+666  Vegetti et al. (2010, 2012); Despali et al. (2024); Ballard et al. (2024); Enzi et al. (2024). Moreover, core collapse alone cannot fully account for the elevated abundance of small-scale lenses reported in galaxy-galaxy strong lensing (GGSL) by Meneghetti et al Meneghetti et al. (2020, 2022, 2023), which exceeds CDM expectations by an order of magnitude. Ref. Dutra et al. (2025) shows that resolving this tension requires extreme central densities in all subhalos, in conflict with the cored profiles needed in dwarfs. Ref. Yang and Yu (2021) further finds that SIDM halos with baryons primarily enhance two-image lensing events, rather than the four-image ones observed.
@@
In this work, we demonstrate that SIDM with mass segregation provides a natural unified framework for resolving these tensions. Mass segregation in multi-component SIDM models arise as a consequence of collisions among particles of different masses. In halos containing two dark matter species with inter-species scattering, energy transfer drives the heavier component toward the center, enhancing central densities Yang et al. (2025a); Patil and Fischer (2025). This process is negligible in CDM due to inefficient gravitational relaxation among particles (unlike stars Zhong and Shapiro (2025)), but becomes significant in SIDM, where collisional interactions dominate. Previous simulations at Milky Way scales Yang et al. (2025a) demonstrated that this mechanism reshapes inner halo profiles under strong scattering. Here, we extend this framework to cosmological simulations of galaxy clusters and controlled simulations of dwarf-scale halos. We show that even modest inter-species interactions can induce mass segregation in massive or high-concentration halos, increasing central densities while leaving dwarf halos largely unaffected. Crucially, unlike in one-component SIDM, the mass enclosed within the Einstein radius can also grow, amplifying strong lensing signals. Our results point to a unified scenario in which both cored dwarf galaxies and cuspy cluster substructures emerge naturally from SIDM with mass segregation.
@@
Figure 1: Effective cross section per mass $\sigma_{\rm eff}/m$ as a function of $V_{\rm max}$ for various SIDM models. The self-interaction among the heavier component ($\chi_{H}-\chi_{H}$) and the inter-species interaction ($\chi_{H}-\chi_{L}$) in the SIDM2v model are represented by red and magenta solid curves. This model is featured by its capability of reproducing SIDM effects analogous to one dark matter component with $\sigma/m=0.3~{}{\rm cm^{2}/g}$ (SIDM03) in massive dwarf halos of mass around $M_{h}=10^{11}~{}M_{\odot}$, as favored by the findings of Ref. Zhang et al. (2025b). The two orange stars feature the two kinds of strengths in the two-component SIDM2c model, which takes constant effective cross sections as in SIDM2v in a $M_{h}=10^{11}~{}\rm M_{\odot}$ halo with median concentration. For $V_{\rm max}\gtrsim 200~{}\mathrm{km/s}$, the inter-species interaction dominates over the intra-species interactions and remains significant up to the cluster scale at $V_{\rm max}\approx 1600~{}\mathrm{km/s}$. In clusters, the coexistence of both types of scattering yields density profiles consistent with observations. Particularly, the intra-species interaction remains well below $0.1~{}\rm cm^{2}/g$, as indicated by the light-red arrow.
@@
A two-component SIDM model with mass segregation— To incorporate mass segregation, we consider a two-component SIDM model in which the heavier species, $\chi_{H}$, has mass $m_{H}=3m_{L}$, with $\chi_{L}$ denoting the lighter component. We assume equal number densities and identical initial spatial distributions for both species. We introduce a model, SIDM2v, featuring velocity-dependent interactions for both identical (intra-species) and distinguishable (inter-species) scatterings. While both cross sections decrease with velocity, the inter-species interaction dominates at high velocities ($V_{\rm max}\gtrsim 200~{}\mathrm{km/s}$). Specifically, we adopt the Møller and Rutherford parametrizations for identical and distinguishable particle scatterings, following equations in Refs. Yang et al. (2025a); Yang and Yu (2022). The intra-species interaction is characterized by $\sigma_{0}/m_{H}=6.89~{}\rm cm^{2}/g$ and a velocity scale $w=275~{}\rm km/s$ (denoted $w_{1}$), while the inter-species interaction uses $\sigma_{0}/m_{H}=1.125~{}\rm cm^{2}/g$ and $w=2200~{}\rm km/s$. We assume that the same particle mediates all intra-species interactions, so that the $\chi_{L}-\chi_{L}$ interaction shares the same $\sigma_{0}$, with a velocity scale $w=(m_{H}/m_{L})w_{1}$.
@@
Figure 1 illustrates our model in terms of effective cross sections, defined as velocity- and angle-averaged differential cross sections weighted by a $v^{5}\sin^{2}\theta$ kernel and convolved with the dark matter velocity distribution. The $x$-axis corresponds to halo velocity scales via their maximum circular velocity, $V_{\rm max}$. The two orange stars indicate the intra- and inter-species interactions in a SIDM2c model without velocity dependence. Their strengths are set by matching the effective cross sections of SIDM2v in a benchmark halo of mass $10^{11}~{}\rm M_{\odot}$, yielding $\sigma_{0}/m=2.25~{}\rm cm^{2}/g$ for intra-species and $\sigma_{0}/m=1.125~{}\rm cm^{2}/g$ for inter-species interactions. This setup leads to time-dependent core formation in most dwarf-hosting halos of mass $\sim 10^{11}~{}\rm M_{\odot}$, considering a single-component SIDM03 model with $\sigma/m=0.3~{}\rm cm^{2}/g$, consistent with the prediction of Ref. Zhang et al. (2025b) and shown as a dashed black line here.
@@
To examine small-scale effects in detail, we perform high-resolution simulations of isolated dwarf halos in SIDM2c, which replicate the relevant interaction regime in SIDM2v. These complement our cosmological simulations near their resolution limit. At higher masses ($M\gtrsim 10^{12}~{}\mathrm{M}_{\odot}$), the intra-species cross section (red curve) drops more steeply than the inter-species one (magenta), leading to enhanced central densities. We also consider a model, SIDMx, that includes only inter-species interactions from SIDM2c, representing an extreme scenario of mass segregation. Notably, both intra- and inter-species interactions are weak yet effective under SIDM2v in clusters. Inter-species interactions drive condensation, which competes with core formation induced by intra-species self-interactions, resulting in a density profile consistent with the $\sigma/m<0.1~{}\rm cm^{2}/g$ constraint from cluster lensing observations Andrade et al. (2021); Kaplinghat et al. (2016); Andrade et al. (2019). We provide an additional supportive test in the Supplemental Material.
@@
Our simulations are based on the SIDM module developed in Refs. Yang and Yu (2022); Yang et al. (2023), extended to handle two dark matter species with unequal masses. This module has been validated in both isolated and cosmological contexts. For isolated halos, we generate NFW profiles in hydrostatic equilibrium using SpherIC Garrison-Kimmel et al. (2013), simulating with $10^{6}$ particles per run. For cosmological simulations, we initialize a periodic box of size $400~{}{\rm Mpc}/h$, and perform zoom-in simulations of a $20~{}{\rm Mpc}/h$ region centered on a host halo of mass $10^{15}~{}{\rm M_{\odot}}/h$, where $h$ is the dimensionless Hubble parameter. The highest-resolution particles have average mass $m_{0}\equiv(m_{1}+m_{2})/2\approx 7\times 10^{7}~{}{\rm M_{\odot}}/h$. As shown in Ref. Yang et al. (2025a), mass segregation is negligible in CDM when particle masses are realistic, so we simulate CDM using a single-component model. For one-component SIDM in isolated halos, we employ the parametric framework of Ref. Yang et al. (2024), which has been extensively tested across controlled and cosmological settings Yang et al. (2025b).
Figure 2: Density profiles of simulated $10^{11}~{}\rm M_{\odot}$ halos with median and ±$\pm 1\sigma$ concentrations at $t=5~{}{\rm Gyr}$ under SIDM03 (dashed) and SIDM2c (solid). The closeness between dashed and solid curves demonstrates that two-component SIDM models retain core formation comparable to the one-component case, consistent with dwarf galaxy clustering observations. The inset further shows that the core formation is time-dependent in two-component models. In the $+1\sigma$ concentration case, we observe that the central densities have clearly increased past their minimum, suggesting an accelerated gravothermal evolution in high concentration halos.

### v2

@@ Abstract
In two-component self-interacting dark matter (SIDM) models with inter-species interactions, mass segregation arises naturally from collisional relaxation, enhancing central densities and gravothermal evolution. We demonstrate that models with velocity-dependent interactions, both within and between species, can connect several small-scale observations while remaining consistent with cluster-scale constraints. This combination enables core formation in dwarf halos, where the presence of baryons increases the inner densities and enhances the predicted strong lensing signatures. Using cosmological and controlled simulations alongside an accurate parametric model, we show that this framework explains the structure of dark perturbers observed in strong lensing systems, and significantly increases the efficiency of small-scale lenses by a factor of $\sim 5$, consistent with the galaxy-galaxy strong lensing excess reported in clusters. Importantly, mass segregation can enhance the Einstein radii of SIDM halos relative to their CDM counterparts, overcoming a key challenge in one-component SIDM scenarios. Our results establish mass segregation in two-component SIDM as a self-consistent and testable model capable of simultaneously addressing multiple small-scale challenges in structure formation.
I Introduction
@@
Observations of large-scale structure have established the standard cosmological model with cold dark matter (CDM). However, on small scales, persistent discrepancies continue to challenge this framework, motivating alternative interpretations Bullock and Boylan-Kolchin (2017); Salucci (2019); Gentile et al. (2004); van Dokkum et al. (2022); Bonaca et al. (2018). Among these, self-interacting dark matter (SIDM) has long been considered a promising candidate Spergel and Steinhardt (2000); Tulin and Yu (2018). Through gravothermal evolution, SIDM can naturally produce both cored and cuspy halo density profiles, providing a flexible mechanism to explain the diversity of galactic rotation curves, the extreme central densities observed in ultra-diffuse and ultra-compact galaxies, and the early formation of supermassive black holes Kaplinghat et al. (2016); Robertson et al. (2019); Despali et al. (2019); Essig et al. (2019); Santos-Santos et al. (2020); Turner et al. (2021); Fischer et al. (2021); Slone et al. (2023); Zeng et al. (2022); Gilman et al. (2021); Correa et al. (2022); Yang and Yu (2022); Adhikari et al. (2022); Outmezguine et al. (2023); Silverman et al. (2022); Yang et al. (2024); Zhong et al. (2023); Zeng et al. (2025); Zhang et al. (2025a, 2024); Ragagnin et al. (2024); Nadler et al. (2025); Kong and Yu (2025); Jiang et al. (2025); Kamionkowski and Sigurdson (2025); Kamionkowski et al. (2025). It has also been proposed as a solution to the final-parsec problem in binary black hole mergers Alonso-Álvarez et al. (2024). Recently, Zhang et al. Zhang et al. (2025b) revealed an unexpected clustering pattern in dwarf galaxies, favoring a tentative self-interaction cross section of $\sigma/m\sim 0.3~\rm cm^{2}/g$ in their host halos.
@@
Despite accumulating evidence, a unified SIDM explanation for all small-scale phenomena remains an open challenge. Observations of dwarf galaxy clustering suggest that a population of these galaxies has growing cores, while the reconstruction of compact substructures inferred from strong lensing images, such as those observed in the SDSS J0946+1006 and JVAS B1938+666 systems Vegetti et al. (2010, 2012); Despali et al. (2024); Ballard et al. (2024); Enzi et al. (2024); Powell et al. (2025), indicates the presence of core-collapsed dwarf halos. Moreover, core collapse alone cannot fully account for the elevated abundance of small-scale lenses reported in galaxy-galaxy strong lensing (GGSL) by Meneghetti et al Meneghetti et al. (2020, 2022, 2023), which exceeds CDM expectations by a factor of 3 to 6. Ref. Yang and Yu (2021) further finds that SIDM halos with baryons primarily enhance two-image lensing events, rather than the four-image ones observed.
@@
In this work, we demonstrate that SIDM with mass segregation provides a natural unified framework for resolving these tensions. Mass segregation in multi-component SIDM models arise as a consequence of collisions among particles of different masses. In the two-component case, inter-species scatterings lead to energy transfer from the heavier to the lighter species, resulting in denser inner halos Yang et al. (2025a); Patil and Fischer (2025). This process is negligible in CDM due to inefficient gravitational relaxation among particles (unlike stars Zhong and Shapiro (2025)), but becomes significant in SIDM, where collisional interactions dominate. Previous simulations at Milky Way scales Yang et al. (2025a) demonstrated that this mechanism reshapes inner halo profiles under strong scattering. Here, we extend this framework to simulations across dwarf to cluster scales, developing a conditioned parametric model to accurately capture gravothermal evolution near the resolution limit. We show that even modest inter-species interactions can induce mass segregation, increasing central densities while producing cores in dwarf halos. Crucially, unlike in one-component SIDM, the mass enclosed within the Einstein radius can also grow, amplifying strong lensing signals. Our results point to a unified scenario in which both cored dwarf galaxies and cuspy cluster substructures emerge naturally from SIDM with mass segregation.
II A two-component SIDM model with mass segregation
@@
Figure 1: Effective cross section per mass $\sigma_{\rm eff}/m$ as a function of $V_{\rm max}$ for various SIDM models. The self-interaction among the heavier component ($\chi_{H}-\chi_{H}$) and the inter-species interaction ($\chi_{H}-\chi_{L}$) in the SIDM2v model are represented by red and magenta solid curves. This model is featured by its capability of reproducing SIDM effects analogous to one dark matter component with $\sigma/m=0.3~{\rm cm^{2}/g}$ (SIDM03) in massive dwarf halos of mass around $M_{h}=10^{11}~\rm M_{\odot}$, as favored by the findings of Ref. Zhang et al. (2025b). The two orange stars feature the two kinds of strengths in the two-component SIDM2c model, which takes constant effective cross sections as in SIDM2v in a $M_{h}=10^{11}~\rm M_{\odot}$ halo with median concentration. For $V_{\rm max}\gtrsim 200~\mathrm{km/s}$, the inter-species interaction dominates over the intra-species interactions and remains significant up to the cluster scale at $V_{\rm max}\approx 1600~\mathrm{km/s}$. In clusters, the coexistence of both types of scattering yields density profiles consistent with observations. Particularly, the intra-species interaction remains well below $0.1~\rm cm^{2}/g$, as indicated by the light-red arrow.
@@
Two-component dark matter models naturally emerge in many particle models, such as those that mirror the asymmetries of the baryonic sector, where protons and electrons are stable and interact with each other Kaplan et al. (2009); Bansal et al. (2023); Petraki et al. (2012); Yang et al. (2025a). To incorporate mass segregation, we consider a two-component SIDM model in which the heavier species, $\chi_{H}$, has mass $m_{H}=3m_{L}$, with $\chi_{L}$ denoting the lighter component. We assume equal number densities and identical initial spatial distributions for both species. We introduce a model, SIDM2v, featuring velocity-dependent interactions for both identical (intra-species) and distinguishable (inter-species) scatterings. While both cross sections decrease with velocity, the inter-species interaction dominates at high velocities ($V_{\rm max}\gtrsim 200~\mathrm{km/s}$). Specifically, we adopt the Møller and Rutherford parametrizations for identical and distinguishable particle scatterings, following equations in Refs. Yang et al. (2025a); Yang and Yu (2022). The intra-species interaction is characterized by $\sigma_{0}/m_{H}=6.89~\rm cm^{2}/g$ and a velocity scale $w=275~\rm km/s$ (denoted $w_{H}$), while the inter-species interaction uses $\sigma_{0}/m_{H}=1.125~\rm cm^{2}/g$ and $w=2200~\rm km/s$. We assume that the same particle mediates all intra-species interactions, so that the $\chi_{L}-\chi_{L}$ interaction shares the same $\sigma_{0}$, with a velocity scale $w=(m_{H}/m_{L})w_{H}$.
@@
Figure 1 illustrates our model in terms of effective cross sections, defined as velocity- and angle-averaged differential cross sections weighted by a $v^{5}\sin^{2}\theta$ kernel and convolved with the dark matter velocity distribution. The $x$-axis corresponds to halo velocity scales via their maximum circular velocity, $V_{\rm max}$. The two orange stars indicate the intra- and inter-species interactions in a SIDM2c model without velocity dependence. Their strengths are set by matching the effective cross sections of SIDM2v in a benchmark halo of mass $10^{11}~\rm M_{\odot}$, yielding $\sigma_{0}/m=2.25~\rm cm^{2}/g$ for intra-species and $\sigma_{0}/m=1.125~\rm cm^{2}/g$ for inter-species interactions. This setup leads to time-dependent core formation in most dwarf halos of mass $\sim 10^{11}~\rm M_{\odot}$, comparable to one-component $\sigma/m=0.3~\rm cm^{2}/g$, as motivated by Ref. Zhang et al. (2025b).
@@
To examine small-scale effects in detail, we perform high-resolution simulations of isolated halos in SIDM2c, which replicate the relevant interaction regime in SIDM2v. We also consider a model, SIDMx, that includes only inter-species interactions from SIDM2c, representing an extreme scenario of mass segregation. Notably, both intra- and inter-species interactions are weak yet effective under SIDM2v in clusters. Inter-species interactions drive condensation, which competes with core formation induced by intra-species self-interactions, resulting in a density profile consistent with the $\sigma/m<0.1~\rm cm^{2}/g$ constraint from cluster lensing observations Andrade et al. (2021); Kaplinghat et al. (2016); Andrade et al. (2019). We provide an additional supportive test in the Supplemental Material. Table 1 summarizes the two-component SIDM models simulated in this work.
Model Scatterings $\frac{\sigma_{0}}{m}$ relation $w$ relation
SIDM2c $\chi_{H}{\texttt{-}}\chi_{H}$ $\frac{\sigma_{H}}{m_{H}}=4.5\rm\frac{cm^{2}}{g}$
$\chi_{L}{\texttt{-}}\chi_{L}$ $\frac{\sigma_{L}}{m_{L}}=\frac{1}{3}\frac{\sigma_{H}}{m_{H}}$
$\chi_{H}{\texttt{-}}\chi_{L}$ $\frac{\sigma_{x}}{m_{H}}=\frac{1}{4}\frac{\sigma_{H}}{m_{H}}$
SIDM2v $\chi_{H}{\texttt{-}}\chi_{H}$ $\frac{\sigma_{H}}{m_{H}}=6.89\rm\frac{cm^{2}}{g}$ $w_{1}=275\ \rm\frac{km}{s}$
$\chi_{L}{\texttt{-}}\chi_{L}$ $\frac{\sigma_{L}}{m_{L}}=\frac{1}{3}\frac{\sigma_{H}}{m_{H}}$ $w_{2}=3w_{1}$
$\chi_{H}{\texttt{-}}\chi_{L}$ $\frac{\sigma_{x}}{m_{H}}=1.125\rm\frac{cm^{2}}{g}$ 2200 $\rm\frac{km}{s}$
SIDMx $\chi_{H}{\texttt{-}}\chi_{L}$ $\frac{\sigma_{x}}{m_{H}}=1.125\rm\frac{cm^{2}}{g}$
Table 1: Benchmark models of two-component SIDM with mass segregation. The $\sigma_{0}/m$ and $w$ parameters are introduced to parametrize the scattering cross sections, following Ref. Yang et al. (2025a); Yang and Yu (2022) and considering Møller (Rutherford) scatterings for identical (distinguishable) particles. For identical particle scattering, a phase space suppression factor of $1/2$ is included to enable comparison with distinguishable inter-species scatterings, yielding the $\sigma_{0}/m=4.5/2{\rm cm^{2}/g}=2.25~\rm cm^{2}/g$ in the SIDM2c case.
III Numerical methods and modeling
To accurately model signatures including baryonic and environmental effects, we employ controlled and cosmological simulations, combined with a parametric model for SIDM2c. The latter is constructed to obtain accurate halo profiles beyond the simulation resolution limit.
@@ III Numerical methods and modeling
Our simulations are based on the SIDM module developed on top of the Gadget-2 program Springel (2005); Springel et al. (2001) in Refs. Yang and Yu (2022); Yang et al. (2023), extended to handle two dark matter species with unequal masses. This module has been validated in both isolated and cosmological contexts. To establish two-component initial conditions without altering the system’s dynamical properties, we randomly select half of the simulation particles to be 50% more massive, and the other half to be 50% less massive, preserving the total system mass. For isolated halos, we generate Navarro-Frenk-White (NFW) profiles Navarro et al. (1997) in hydrostatic equilibrium using SpherIC Garrison-Kimmel et al. (2013), simulating with $N=10^{6}$ particles per run. The softening length is chosen to be a few $R_{\rm vir}/\sqrt{N}$ Power et al. (2003), where $R_{\rm vir}$ is the virial radius. A stellar component, modeled by the Hernquist profile Hernquist (1990), is included in the study with baryons. They are simulated as collisionless particles with the same mass as the dark matter particles. In the Supplemental Material, we provide a summary of the controlled simulations in a table. For cosmological simulations, we initialize a periodic box of size $400~{\rm Mpc}/h$, and perform zoom-in simulations of a $20~{\rm Mpc}/h$ region centered on a host halo of mass $10^{15}~{\rm M_{\odot}}/h$, with a softening length of one comoving $kpc/h$, where $h$ is the dimensionless Hubble parameter. The highest-resolution particles have average mass $m_{0}\equiv(m_{1}+m_{2})/2\approx 7\times 10^{7}~{\rm M_{\odot}}/h$. As shown in Ref. Yang et al. (2025a), mass segregation is negligible in CDM when particle masses are realistic, so we simulate CDM using a single-component model. We also apply a modified version of the Rockstar halo finder Behroozi et al. (2013) to accommodate two-component dark matter, a method validated and applied in Ref. Yang et al. (2025a). Using the resulting halo catalogs, we then build halo merger trees with the consistent-trees program Behroozi et al. (2013). For one-component SIDM in isolated halos, we employ the parametric framework of Ref. Yang et al. (2024), which has been extensively tested across controlled and cosmological settings Yang et al. (2025b).
To construct a parametric model for SIDM with mass segregation, we first show that isolated halos in our setup exhibit conditioned universality. Within the conduction-fluid framework, the coupled transfer equations for the heavy ($H$) and light ($L$) dark matter species are
$\rho_{H}\nu_{H}^{2}\left(\frac{D}{Dt}\right)\ln\frac{\nu_{H}^{3}}{\rho_{H}}$ $=$ $-\frac{1}{4\pi r^{2}}\frac{\partial L_{H}}{\partial r}-R,$ (1)
$\rho_{H}\nu_{H}^{2}\left(\frac{D}{Dt}\right)\ln\frac{\nu_{L}^{3}}{\rho_{L}}$ $=$ $-\frac{1}{4\pi r^{2}}\frac{\partial L_{L}}{\partial r}+R,$
where $\rho,\nu$ and $L$ denote density, velocity dispersion, and luminosity, respectively. Here, $D/Dt$ is the material derivative and $R$ the inter-species energy transfer rate.
Figure 2: Rescaled halo density profiles at 60% of the core collapse time $\tau\equiv t/t_{c}=0.6$, compared to a single-component (SIDM1c) model at the same evolution stage. The profiles of three halos of mass $2\times 10^{10}~M_{\odot}$ (solid) and three of mass $10^{11}~M_{\odot}$ (dashed) with median, $1\sigma$, and $1\sigma$ concentrations have almost identical shapes, illustrating the high quality of universality supporting the construction of the parametric model.
Figure 3: Dwarf halo evolution and core formation in one- and two-component SIDM. Left: Density profiles for the median halo with a collisionless stellar component at $t=0$, 33, 66, and 99~Gyr, with deeper colors corresponding to later times. Baryons (dotted tomato lines) induce higher DM densities (solid) compared to DM-only cases (dashed). Unlike DM-only evolution, central density increases even as the cores grow in size. The parametric model predictions are shown in dash-dotted lines. Right: Evolution of core size ($r_{c}$) for the median halo with (solid and dashed) and without (dotted) baryons. The core growth in SIDM2c remains comparable to that in SIDM03 in the presence of baryons.
In the long-mean-free-path regime, the SIDM effect scales with the scattering rate, hence $L_{H},L_{R}$ and $R$ are all proportional to the scattering cross section. In two-component models, the normalization of these quantities, parameterized by $\sigma_{s}/m$, can be absorbed into the evolution time, defining a rescaled evolution variable $t\sigma_{s}/m$. It follows that SIDM halos exhibit a universal evolutionary track when the mass ratio and the relative inter- to intra-species cross sections are fixed.
Figure 2 illustrates the accuracy of this conditioned universality by showing that the rescaled halo density profiles have almost the same shapes if evolved under SIDM2c for 60% of their core collapse times. We consider two halos of masses $2\times 10^{10}~\rm M_{\odot}$ and $10^{11}~\rm M_{\odot}$, with varied concentrations computed following the $c_{200}-M$ relation from Ref. Dutton and Macciò (2014) and assuming a logarithmic scatter of $0.11$ dex. Compared with the density profile of a halo in one-component SIDM (SIDM 1c) at the same gravothermal phase ($\tau\equiv t/t_{c}=0.6$, shown in dotted, tomato-colored curve), the profiles in SIDM2c have cored profiles and significantly higher inner densities.
By fitting the universal density profile evolution, we construct an accurate parametric model for halos under SIDM2c. We also extend the model to capture leading baryon effect on the profile evolution, which is achieved by simply rescaling the core size in the DM-only case by a form factor introduced in Ref. Yang (2024). In Ref. Meneghetti et al. (2022); Bahé (2021); Robertson (2021); Tokayer et al. (2024), it has been demonstrated that baryons have a significant effect on cluster subhalos. While simplified, our method captures baryons’ contraction on SIDM2c evolution, the effect of stellar-to-halo mass ratio, and reduces to the DM-only case without baryons. We provide supplemental material to detail the model and demonstrate its robustness.
Figure 4: Distribution of the gravothermal phase $\tau$ from parametric model predictions for the SIDM2c model. The sample includes halos of masses higher than $10^{10}~\rm M_{\odot}$ and within $5~\mathrm{Mpc}/h$ of the host cluster, separated into cluster subhalos (Group A, light red) within the cluster’s virial radius and isolated halos (Group B, light blue). Means and standard deviations for both groups are shown in the inset. Most halos in both groups have cored inner profiles ($\tau<0.6$). Cluster subhalos, however, show a greater frequency of core collapsed systems ($\tau>1$), as a consequence of tidal stripping.
IV Dwarf scale cores